OpenStudy (anonymous):

(b^4)^6 * (b^2)^4 Simplify ? i need help

OpenStudy (kirbykirby):

(b^4)^6=b^24 (b^2)^4=b^8 b^24*b^8 = b^32 When you have the form (a^x)^y... u multiply the exponents so a^(xy) When you have the form b^x*b^y, you add the exponents so b^(x+y)

OpenStudy (anonymous):

ok so does the formula stay the same if the problem read (k^3d^5)^5 ? im getting a bit confused on this one

OpenStudy (kirbykirby):

yes you "distribute" the ^5... so you get k^15 * d^25

OpenStudy (kirbykirby):

you can't simplify further since you have 2 different variables

OpenStudy (anonymous):

thanks so much :) could you help me with a area and dimensions problem?...i dont want the answer i just want to know how to set up the problem

OpenStudy (kirbykirby):

I could try my best :)

OpenStudy (anonymous):

Justin wants to use 188 ft of fencing to fence off the greatest possible rectangular area for a garden. how would i find what dimensions he should use and what the area would be

OpenStudy (kirbykirby):

Is this a calculus question

OpenStudy (anonymous):

yea i think so

OpenStudy (kirbykirby):

Ok so you want to maximize the area of the rectangle. The area of a rectangle is A=b*L (area=base*length). We know the perimeter of a rectangle is 2(b+L)=188. Since you want to maximize A, you want to express A with only one variable. So, you use the perimeter equation above and solve, for say L, and substitute the expression for L into your Area formula. So now you have an equation A = ... (which only has one variable "b"). So, know you just differentiate your function and setting it equal to 0 to find the critical points. Make sure you only consider actual physical values (if you get a negative value, for example, ignore it since you can't have a negative length). So, then you evaluate your function A = ... at the critical points and determine which value gives the largest value for A. This will be your maximum area.

OpenStudy (kirbykirby):

I hope that's clear :S

OpenStudy (kirbykirby):

It's funny you ask this question because I have a review assignment for another class based on first year calculus and the problem I am attacking now is an optimization problem as well :P

OpenStudy (kirbykirby):

Oh wait I partially answered the question... You are asking for the dimensions. So, the process I told you will let you find "b". So with that, you can plug is the value for b in the perimeter equation 2(b+L)=188 and solve for L to get the length

OpenStudy (anonymous):

its very clear just hard to understand because i have ADD which makes it kind of hard to read everything without forgetting -_- im going to test it and post my answer can you tell me if im right?

OpenStudy (kirbykirby):

Ok sure

OpenStudy (anonymous):

i got 8,811 ft^2 ?

OpenStudy (kirbykirby):

Um I got 2209

OpenStudy (anonymous):

:/ i went wrong somewhere

OpenStudy (kirbykirby):

If you think of it in a simple way, although this depends if your prof talked about it or not, but the greatest area rectangle you can get is always a square. So 188 = 4c (c = side of a square) so c = 47. So each side is 47 ft. The area of a square = c^2 = 47^2 = 2209

OpenStudy (kirbykirby):

Do you know how u got 8811?

OpenStudy (anonymous):

thats what i just got 47X47 47; 2,209 ft2 :) i went back over your steps..maths so hard !

OpenStudy (kirbykirby):

Yay awesome :)!

OpenStudy (anonymous):

one more favor ...can you help me simplify an expression? i pretty much got the rest

OpenStudy (kirbykirby):

Ok sure

OpenStudy (anonymous):

4x^2y^7 over 24x^3y^4

OpenStudy (kirbykirby):

\[\frac{ 4x^2y^7 }{ 24x^3y^4 }=\frac{ 1x^{2-3}*y^{7-4} }{ 6 }=\frac{ x^{-1}*y^{3} }{ 6 }=\frac{ y^{3} }{ 6x }\]

OpenStudy (anonymous):

THANKS SO MUCH! wish i could give you a million medals :)

OpenStudy (kirbykirby):

hm ok the numbers look small but it says x^(2-3) in the second "="

OpenStudy (kirbykirby):

:) No problem!